40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Feb 5 02:04:39 2015, Zach, DailyProgress, SiFi - ringdown, No big Q increase at low temperature 
    Reply  Fri Feb 6 04:23:20 2015, Zach, DailyProgress, SiFi - ringdown, No big Q increase at low temperature 
Message ID: 1194     Entry time: Fri Feb 6 04:23:20 2015     In reply to: 1193
Author: Zach 
Type: DailyProgress 
Category: SiFi - ringdown 
Subject: No big Q increase at low temperature 

I monitored the reservoir level periodically over the day and night. As of the evening, there appeared to be ~1 cm of LN2 still there. As of around 4am, it appears empty, so it should be OK to open tomorrow. I've sealed the vacuum and shut off the pump in preparation.

Quote:

I'm going to wait for things to warm up and then vent the chamber so that we can:

  1. Improve the clamp
  2. Fix our wiring issues

 

ELOG V3.1.3-